# LDPRecover: Recovering Frequencies from Poisoning Attacks against Local Differential Privacy

### v1

@@ III-A Local Differential Privacy
A randomized algorithm Ψ()normal-$\Psi(\cdot)$ satisfies $\epsilon$-LDP, where $\epsilon>=0$, if and only if for any two input $v_{1},v_{2}\in D$, we have
@@ V-C Malicious Frequency Learning
$\mathbf{E}[\sum\nolimits_{v\in D}\tilde{f}_{Y}(v)]=\frac{\sum_{v\in D}P(v)-qd}{p-q}=\frac{1-qd}{p-q}$ (20)
where $P(v)$ is the probability of item $v$ in $P$, $d$ is the size of $D$, and $p,q$ are perturbation probabilities of the LDP protocol. Note that the value of $\mathbf{E}[\sum_{v\in D}\tilde{f}_{Y}(v)]$ changes with varying parameters $p$ and $q$ in different LDP protocols. This is due to the fact that the data reported by malicious users bypasses the perturbation algorithm of the LDP protocol, but is subjected to the aggregation algorithm of the LDP protocol. Additionally, regardless of the attacker-designed distribution $P$, the sum of frequencies across all items always be $1$, i.e., $\sum_{v\in D}P(v)=1$. Consequently, we can use the expected summation to approximate the actual summation of malicious frequencies for all items, i.e.,
@@ VI-A4 Recovery Settings
For ease of presentation, we use LDPRecover and $*{}^{*}$ to denote the versions of LDPRecover operating without and with partial knowledge of the items selected by the attacker, respectively. In LDPRecover, the server aggregates item frequencies without any detail of the attacks. Conversely, $*{}^{*}$ operates under the assumption that the server is aware of the attacker-selected items. In the context of MGA, these items are explicitly identified as target items, while in AA, they are the items that exhibit the top-$r/2$ frequency increase following the attack.
The parameter involved in LDPRecover and $*{}^{*}$ is $\eta=\frac{m}{n}$ (the ratio of the number of malicious users to the number of genuine users). Since, in practice, the server does not know what the real value of $\eta$ is, we set a large $\eta=0.2$ by default in the experiments, which is a value larger than the real value (i.e., $\eta\gg\frac{\beta}{1-\beta}$). We also explore the impact of $\eta$ on recovered results from the poisoning attacks in Section VI-D, where the range of $\eta$ is $\eta\in[0.01,0.4]$.
@@ VI-A5 Comparison Methods
Figure 3: The mean squared error (MSE) of LDPRecover and $*{}^{*}$ for two datasets, three LDP protocols, and three attacks. “-Manip”, “-MGA”, and “-AA” represents the results for recovery from Manip, MGA, and AA, respectively.
Figure 4: The frequency gain (FG) of LDPRecover and $*{}^{*}$ for two datasets, three LDP protocols, and three attacks. “-Manip”, “-MGA”, and “-AA” represents the results for recovery from Manip, MGA, and AA, respectively.
@@ VI-C Experimental Results
Figures 3 and 4 show the MSE and FG of LDPRecover and $*{}^{*}$ for the two datasets, three LDP protocols, and three attacks. From these figures, we have the following observations:
Both LDPRecover and $*{}^{*}$ are Accurate and Widely Applicable. Both LDPRecover and $*{}^{*}$ can recover the accurate aggregated frequencies from the poisoned ones caused by Manip, MGA, and AA, even if the default $\eta=0.2$ significantly exceeds the actual ratio $\frac{\beta}{1-\beta}=0.052$. Notably, $*{}^{*}$ consistently performs best when countering MGA, i.e., it achieves a lower MSE than LDPRecover. Specifically, both LDPRecover and $*{}^{*}$ recover aggregated frequencies by subtracting the malicious frequencies from the poisoned frequencies, i.e., the more accurate the estimated malicious frequencies, the more accurate the recovered frequencies are. Based on this, while LDPRecover estimates these malicious frequencies based solely on the LDP protocol’s properties, $*{}^{*}$ refines these estimations by incorporating information about the attacker-selected items, thus enhancing accuracy. To confirm this, we evaluated the MSEs of the malicious frequencies estimated by LDPRecover and $*{}^{*}$ versus the true malicious frequencies. The experimental results in Figure 7 show that $*{}^{*}$ estimates malicious frequencies more accurately than LDPRecover. Consequently, $*{}^{*}$ surpasses LDPRecover in recovering the aggregated frequencies by estimating more accurate malicious frequencies. In other words, the prior knowledge of attacker-selected items introduced in $*{}^{*}$ helps to improve the accuracy of the recovered frequency. In addition, we note that LDPRecover and $*{}^{*}$ outperform Detection in all cases. This is because Detection removes all users with the target items indiscriminately, such that the genuine users with the target items are incorrectly removed.
Figure 7: MSE between LDPRecover and $*{}^{*}$ estimated malicious frequencies and true malicious frequencies (IPUMS).
Both LDPRecover and $*{}^{*}$ reduce FG of Targeted Attacks. LDPRecover and $*{}^{*}$ can effectively defend against targeted attacks. First, LDPRecover can significantly reduce the FG of targeted attacks, especially to almost 00 in most cases. Second, $*{}^{*}$ even reduces the FG to a negative value (FG$<0$), which means that the recovered frequencies of the target items are even small than the ones before the attack. Third, Detection is much less effective in dealing with targeted attacks than LDPRecover, not to mention $*{}^{*}$. This performance gap comes from the same reason: Detection brutally removes all users with target items. In contrast, $*{}^{*}$, which has the same prior knowledge as Detection, can obtain more accurate malicious data, as shown in Figure 7. This allows $*{}^{*}$ to finely deduct malicious data from poisoned data, gaining a lower FG.
@@ VI-D Effectiveness of Various Parameters
Impact of the fraction of malicious users $\beta$. The first column of Figures 5 and 6 shows the impact of $\beta$ on recovering frequencies from AA over IPUMS and Fire datasets, respectively. We observe that when defensing against poisoning attacks, LDPRecover and $*{}^{*}$ can recover the accurate aggregated frequencies under various $\beta$. Moreover, Table I shows the MSE of the LDPRecover execution on the unpoisoned frequencies, where Before-Rec and After-Rec denote the MSE of unpoisoned frequencies and the MSE of the frequencies recovered by LDPRecover on the unpoisoned frequencies, respectively. From this table, we observe that when LDPRecover executes on the unpoisoned frequencies, it can improve the accuracy of the frequencies from GRR and reduce the accuracy of the frequencies from OLH and OLH. This is because when LDPRecover executes on the unpoisoned frequencies from OUE and OLH, it may remove frequencies that should not be removed, thus reducing accuracy.
Impact of the privacy budget of the LDP protocols $\epsilon$. The second column of Figures 5 and 6 shows the impact of the privacy budget on recovering frequencies from AA over IPUMS and Fire datasets, respectively. We observe that regardless of $\epsilon$, LDPRecover and $*{}^{*}$ are both effective in recovering accurate aggregated frequencies. In particular, for $*{}^{*}$, the MSE remains low and stable under in all cases; for LDPRecover, the MSE may decreases or remains stable as $\epsilon$ grows. The discrepancy arises because $*{}^{*}$ exploits known details of the poisoning attack, resulting in a stable and accurate estimation. In contrast, LDPRecover utilizes the LDP protocol’s properties for estimation, rendering it susceptible to fluctuations caused by $\epsilon$.
Impact of the ratio of the number of malicious users to the number of genuine users $\eta$. The third column of Figures 5 and 6 shows the effect of $\eta$ on LDPRecover and $*{}^{*}$. We observe that LDPRecover and $*{}^{*}$ perform optimally when $\eta$ closely matches $\beta$. This is because a more accurate $\eta$ will enable our methods to estimate genuine frequencies more accurately. Furthermore, they still maintain effectiveness even there is a deviation between $\eta$ and $\beta$. This is because, even if the estimated genuine frequencies are moderately accurate, our method can still refine accurate recovered frequencies from the estimated genuine frequencies using public available knowledge (i.e., Equations (22) and (23)). For example, Figure 5 (a) illustrates that with $\beta=0.05$ and $\eta=0.4$, LDPRecover significantly outperforms the poisoned data, with the average MSE of $1.42\times 10^{-4}$ for LDPRecover versus $8.78\times 10^{-2}$ for the poisoned frequencies.

### v2

@@ III-A Local Differential Privacy
A randomized algorithm $\Psi(\cdot)$ satisfies $\epsilon$-LDP, where $\epsilon>=0$, if and only if for any two input $v_{1},v_{2}\in D$, we have
@@ V-C Malicious Frequency Learning
$\mathbf{E}[\sum\nolimits_{v\in D}\tilde{f}_{Y}(v)]=\frac{\sum_{v\in D}f_{Y}(v)-qd}{p-q}=\frac{1-qd}{p-q}$ (20)
where $f_{Y}(v)$ is the frequency of item $v$ in $Y$, $d$ is the size of $D$, and $p,q$ are perturbation probabilities of the LDP protocol. Note that the value of $\mathbf{E}[\sum_{v\in D}\tilde{f}_{Y}(v)]$ changes with varying parameters $p$ and $q$ in different LDP protocols. This is due to the fact that the data reported by malicious users bypasses the perturbation algorithm of the LDP protocol, but is subjected to the aggregation algorithm of the LDP protocol. Additionally, regardless of the attacker-designed distribution $P$, we always have $\sum_{v\in D}f_{Y}(v)=1$. Consequently, we can use the expected summation to approximate the actual summation of malicious frequencies for all items, i.e.,
@@ VI-A4 Recovery Settings
For ease of presentation, we use LDPRecover and LDPRecover to denote the versions of LDPRecover operating without and with partial knowledge of the items selected by the attacker, respectively. In LDPRecover, the server aggregates item frequencies without any detail of the attacks. Conversely, LDPRecover operates under the assumption that the server is aware of the attacker-selected items. In the context of MGA, these items are explicitly identified as target items, while in AA, they are the items that exhibit the top-$r/2$ frequency increase following the attack.
The parameter involved in LDPRecover and LDPRecover is $\eta=\frac{m}{n}$ (the ratio of the number of malicious users to the number of genuine users). Since, in practice, the server does not know what the real value of $\eta$ is, we set a large $\eta=0.2$ by default in the experiments, which is a value larger than the real value (i.e., $\eta\gg\frac{\beta}{1-\beta}$). We also explore the impact of $\eta$ on recovered results from the poisoning attacks in Section VI-D, where the range of $\eta$ is $\eta\in[0.01,0.4]$.
@@ VI-A5 Comparison Methods
Figure 3: The mean squared error (MSE) of LDPRecover and LDPRecover for two datasets, three LDP protocols, and three attacks. “-Manip”, “-MGA”, and “-AA” represents the results for recovery from Manip, MGA, and AA, respectively.
Figure 4: The frequency gain (FG) of LDPRecover and LDPRecover for two datasets, three LDP protocols, and three attacks. “-Manip”, “-MGA”, and “-AA” represents the results for recovery from Manip, MGA, and AA, respectively.
@@ VI-C Experimental Results
Figures 3 and 4 show the MSE and FG of LDPRecover and LDPRecover for the two datasets, three LDP protocols, and three attacks. From these figures, we have the following observations:
Both LDPRecover and LDPRecover are Accurate and Widely Applicable. Both LDPRecover and LDPRecover can recover the accurate aggregated frequencies from the poisoned ones caused by Manip, MGA, and AA, even if the default $\eta=0.2$ significantly exceeds the actual ratio $\frac{\beta}{1-\beta}=0.052$. Notably, LDPRecover consistently performs best when countering MGA, i.e., it achieves a lower MSE than LDPRecover. Specifically, both LDPRecover and LDPRecover recover aggregated frequencies by subtracting the malicious frequencies from the poisoned frequencies, i.e., the more accurate the estimated malicious frequencies, the more accurate the recovered frequencies are. Based on this, while LDPRecover estimates these malicious frequencies based solely on the LDP protocol’s properties, LDPRecover refines these estimations by incorporating information about the attacker-selected items, thus enhancing accuracy. To confirm this, we evaluated the MSEs of the malicious frequencies estimated by LDPRecover and LDPRecover versus the true malicious frequencies. The experimental results in Figure 7 show that LDPRecover estimates malicious frequencies more accurately than LDPRecover. Consequently, LDPRecover surpasses LDPRecover in recovering the aggregated frequencies by estimating more accurate malicious frequencies. In other words, the prior knowledge of attacker-selected items introduced in LDPRecover helps to improve the accuracy of the recovered frequency. In addition, we note that LDPRecover and LDPRecover outperform Detection in all cases. This is because Detection removes all users with the target items indiscriminately, such that the genuine users with the target items are incorrectly removed.
Figure 7: MSE between LDPRecover and LDPRecover estimated malicious frequencies and true malicious frequencies (IPUMS).
Both LDPRecover and LDPRecover reduce FG of Targeted Attacks. LDPRecover and LDPRecover can effectively defend against targeted attacks. First, LDPRecover can significantly reduce the FG of targeted attacks, especially to almost 00 in most cases. Second, LDPRecover even reduces the FG to a negative value (FG$<0$), which means that the recovered frequencies of the target items are even small than the ones before the attack. Third, Detection is much less effective in dealing with targeted attacks than LDPRecover, not to mention LDPRecover. This performance gap comes from the same reason: Detection brutally removes all users with target items. In contrast, LDPRecover, which has the same prior knowledge as Detection, can obtain more accurate malicious data, as shown in Figure 7. This allows LDPRecover to finely deduct malicious data from poisoned data, gaining a lower FG.
@@ VI-D Effectiveness of Various Parameters
Impact of the fraction of malicious users $\beta$. The first column of Figures 5 and 6 shows the impact of $\beta$ on recovering frequencies from AA over IPUMS and Fire datasets, respectively. We observe that when defensing against poisoning attacks, LDPRecover and LDPRecover can recover the accurate aggregated frequencies under various $\beta$. Moreover, Table I shows the MSE of the LDPRecover execution on the unpoisoned frequencies, where Before-Rec and After-Rec denote the MSE of unpoisoned frequencies and the MSE of the frequencies recovered by LDPRecover on the unpoisoned frequencies, respectively. From this table, we observe that when LDPRecover executes on the unpoisoned frequencies, it can improve the accuracy of the frequencies from GRR and reduce the accuracy of the frequencies from OLH and OLH. This is because when LDPRecover executes on the unpoisoned frequencies from OUE and OLH, it may remove frequencies that should not be removed, thus reducing accuracy.
Impact of the privacy budget of the LDP protocols $\epsilon$. The second column of Figures 5 and 6 shows the impact of the privacy budget on recovering frequencies from AA over IPUMS and Fire datasets, respectively. We observe that regardless of $\epsilon$, LDPRecover and LDPRecover are both effective in recovering accurate aggregated frequencies. In particular, for LDPRecover, the MSE remains low and stable under in all cases; for LDPRecover, the MSE may decreases or remains stable as $\epsilon$ grows. The discrepancy arises because LDPRecover exploits known details of the poisoning attack, resulting in a stable and accurate estimation. In contrast, LDPRecover utilizes the LDP protocol’s properties for estimation, rendering it susceptible to fluctuations caused by $\epsilon$.
Impact of the ratio of the number of malicious users to the number of genuine users $\eta$. The third column of Figures 5 and 6 shows the effect of $\eta$ on LDPRecover and LDPRecover. We observe that LDPRecover and LDPRecover perform optimally when $\eta$ closely matches $\beta$. This is because a more accurate $\eta$ will enable our methods to estimate genuine frequencies more accurately. Furthermore, they still maintain effectiveness even there is a deviation between $\eta$ and $\beta$. This is because, even if the estimated genuine frequencies are moderately accurate, our method can still refine accurate recovered frequencies from the estimated genuine frequencies using public available knowledge (i.e., Equations (22) and (23)). For example, Figure 5 (a) illustrates that with $\beta=0.05$ and $\eta=0.4$, LDPRecover significantly outperforms the poisoned data, with the average MSE of $1.42\times 10^{-4}$ for LDPRecover versus $8.78\times 10^{-2}$ for the poisoned frequencies.